OpenStudy (anonymous):

Newtons recipes for 32 biscuits calls for 2 and 1/4 of flour. How much flour will Newton need to make 224 biscuits? F. 15 cups G. 15 and 1/2 cups H. 15 and 3/4 cups J. 18 and 1/2 cups

OpenStudy (anonymous):

convert 2 1/4 into proper function =9/4 32 biscuits require 2 and 1/4 then 224=32*x=>(9/4)*x here u have to find x and substitute in (9/4)*x u will get answer

OpenStudy (anonymous):

I don't get you o.o

OpenStudy (anonymous):

i mean 2 1/4 should be (2*4+1)/4, u get 9/4 u know that 224=32*8 if 32 biscuits take 9/4 flour, then 32*7 takes= 9/4*7=63/4 flour convert it back to improper function

OpenStudy (anonymous):

or u can directly multiply, 2 and 1/4 with 7, u get 14 and 7/4, and 7/4 can be written as 1 and 3/4 so 15 3/4 should be ans

OpenStudy (anonymous):

Thank you